1.     A company manufactures soaps. The mean mass of the soaps is 300g with variance 52g. A sample of 100soaps was selected.

a)     Find the probability of selecting soaps more than masses 320g.

b)     How many soaps will have masses between 280g and 500g.

 


1
Expert's answer
2022-03-01T15:47:18-0500

Let "X=" mass of the soap: "X\sim N(\mu, \sigma^2)."

Given "\mu=300\ g, \sigma=\sqrt{52}\ g."

a)


"P(X>320)=1-P(Z\le\dfrac{320-300}{\sqrt{52}})"

"\approx1-P(Z\le2.7735)\approx0.002773"


b)


"P(280<X<500)=P(Z<\dfrac{500-300}{\sqrt{52}})"

"-P(Z\le\dfrac{280-300}{\sqrt{52}})\approx P(Z\le27.7350)"

"-P(Z\le-2.7735)\approx1-0.002773"

"\approx0.997227"

"0.997227(100)=100"

100 soaps will have masses between 280g and 500g.